OpenStudy (anonymous):

Use newton's method of approximation to approximate the zero of the function in the indicated interval to six decimal places f(x) = cos x -x in [0,pi/2] Our working has to be in radians...but i'm not quite sure how to do that... i know for my x1 it can be pi/4

OpenStudy (anonymous):

i would like to know how to do it manually...not on the computer :) thank you :)

myininaya (myininaya):

f'(x)=-sinx-1 we need to choose an x to start with like x=1/2 \[x_n=x_{n-1}-\frac{f(x_{n-1})}{f'(x_{n-1}}\] so x_1=1/2 x_2=1/2-f(1/2)/f'(1/2)=.692749953 then you find x_3 using same process i do this until my x's start repeating

OpenStudy (anonymous):

i think it has to be in radians so i thought we use x=pi/4

myininaya (myininaya):

x_3=.692749953-f(.692749953)/f'(.692749953)=.7395826723

myininaya (myininaya):

it is in radians

OpenStudy (anonymous):

oh my bad so why u choosed x = 1/2 because the interval is continuous at that point?

myininaya (myininaya):

i chose 1/2 because it was closer to the x-intercept of the curve you can do pi/4 if you want

myininaya (myininaya):

7/10 probably would have been closer

OpenStudy (anonymous):

hmm okay let me see how that works out

OpenStudy (anonymous):

i thought were focusing on the interval [0,pi/2] and we want a number between that :s lol

myininaya (myininaya):

we are using the x-intercepts of tangent lines to approximate the x-intercept of the curve

OpenStudy (anonymous):

hmm...:S

myininaya (myininaya):

.5 is between 0 and pi/2

myininaya (myininaya):

pi/2 is approximately 1.570796 blah blah

OpenStudy (anonymous):

pi/2 is .5 no...?

OpenStudy (anonymous):

lol...

myininaya (myininaya):

pi is approximately 3.14 blah blah 3.14/2 is bigger than one so .5 is definitely not 1

OpenStudy (anonymous):

hmm sorry i'm forgetting my trig..how do you convert from degrees to radians... out of your head ?lol

myininaya (myininaya):

30 degrees=pi/6 right? so the question is how do we get from 30 degrees to pi/6 so 30*pi/180 so you just multiply the number in degrees by pi/180 (canceling the degrees on both top and bottom since 180 is in degrees also)

myininaya (myininaya):

to get from pi/6 to 30 degrees you multiply pi/6 by 180 degrees/pi

OpenStudy (anonymous):

ahh yes ok thank you...it's all coming back now

myininaya (myininaya):

radians is the same as if we are talk about 1 (it can be integer) if you want if it makes since you can call your radians normal numbers lol does that make since you aren't working with degrees but normal numbers

OpenStudy (anonymous):

lol okay

myininaya (myininaya):

do you have anymore questions

myininaya (myininaya):

do you have i ti83?

OpenStudy (anonymous):

umm just trying to do it using pi/4...but i think i get the hang of it

OpenStudy (anonymous):

no...i have a casio fx-350MS is that a graphing calc

myininaya (myininaya):

go to mode and make sure your calculator is in radians

OpenStudy (anonymous):

yes i did that thank you

myininaya (myininaya):

ok it looks like you are all set cool!

OpenStudy (anonymous):

lol thank so much again

myininaya (myininaya):

:)